# nGPT: Normalized Transformer with Representation Learning on the Hypersphere

### v1

@@ 1 Introduction
We propose to unify various findings and observations made in the field under a new perspective of the normalized Transformer. Our key contributions are as follows:
@@
Faster convergence We demonstrate that the normalized Transformer reduces the number of training steps required to achieve the same accuracy by a factor of 4 to 20.
@@ 2.1 Token Embeddings and Output Logits
Here, $z_{i,y_{i}}$ is the logit corresponding to the correct token $y_{i}$, and the denominator normalizes the logits into a probability distribution over the vocabulary. During inference, the prediction $\hat{{\bm{y}}}_{i}$ is obtained by selecting the token with the highest probability. Since all nGPT embeddings are normalized, the logits ${\bm{z}}\in\mathbb{R}^{V}$ in equation 1 represent dot products bounded in the range $[-1,1]$. This limits the confidence (temperature) of the probability distribution generated by the softmax in equation 2. To adjust this during training, we introduce a trainable scaling parameter ${\bm{s}}_{z}\in\mathbb{R}^{V}$ that scales the logits element-wise:
@@ 2.6 Summary of modifications
    After each training step, normalize all matrices (${\bm{E}}_{\text{input}}$, ${\bm{E}}_{\text{output}}$, ${\bm{W}}_{q}$, ${\bm{W}}_{k}$, ${\bm{W}}_{v}$, ${\bm{W}}_{o}$, ${\bm{W}}_{u}$, ${\bm{W}}_{\nu}$ and ${\bm{W}}_{o\text{MLP}}$) along their embedding dimension.
@@ 3.1 Acceleration of training
Figure 3 shows a similar pattern across downstream tasks, confirming that the acceleration is not only reflected in perplexity but also in task performance. Figures 8 and 9 in the Appendix provide results for 1k and 8k context lengths. We observe some saturation for the longest runs of nGPT, suggesting that the model capacity is nearly reached for this number of trainable model parameters.
@@ 3.2 Inspection of network parameters
An important contribution of this work is the decoupling of predictions made by the Attention and MLP blocks from their impact on the hidden state ${\bm{h}}$. These contributions are controlled by the eigen learning rates $\bm{\alpha}_{\text{A}}$ and $\bm{\alpha}_{\text{M}}$. Their interpretation is straightforward: if $\bm{\alpha}_{\text{A},i}$ for an embedding dimension $i\in\mathbb{R}^{d_{\text{model}}}$ is 0.2, then the update follows $h_{i}\leftarrow(1-0.2)h_{i}+0.2h_{\text{A},i}$. Thus, they directly quantify the contribution of $h_{\text{A},i}$ into ${\bm{h}}$. Figure 6 shows the average absolute values of ${\bm{h}}_{\text{A}}$ and ${\bm{h}}_{\text{M}}$ at each layer. Notably, the network learns to take only modest steps (20%-30%) in the direction suggested by ${\bm{h}}_{\text{A}}$ and ${\bm{h}}_{\text{M}}$. The average magnitude of $\bm{\alpha}_{\text{A}}$ decreases from 0.25 in the 0.5B network (24 layers) to 0.20 in the 1B network (36 layers). Meanwhile, $\bm{\alpha}_{\text{M}}$ decreases from 0.37 to 0.32, possibly because MLP blocks have more parameters, making their suggestions more precise.
@@ 4 Related Work
The normalization in equations 15 and 16 closely resembles the QK normalization by Henry et al. (2020). In nGPT, this process can be viewed as restoring ${\bm{q}}$ and ${\bm{k}}$ of the $i$-th head to a $(d_{\text{model}}/n_{\text{heads}})$-dimensional hypersphere after the projection of ${\bm{h}}$ by ${\bm{W}}_{q}$ and ${\bm{W}}_{k}$, respectively. Since ${\bm{h}}$ and the embedding dimensions of ${\bm{W}}_{q}$ and ${\bm{W}}_{k}$ are already normalized, the norms of ${\bm{q}}$ and ${\bm{k}}$ are also comparable, making their normalization potentially unnecessary. We explored this scenario (i.e., omitting the normalization of ${\bm{q}}$ and ${\bm{k}}$) in our ablation studies (see Appendix A.8), where the results showed only a slight degradation. The performance drop was comparable to the computational time savings per step. Therefore, removing the normalization of ${\bm{q}}$ and ${\bm{k}}$ in nGPT is a viable option.
@@ 5 Discussion and Conclusion
This work builds on numerous key findings and observations made in the field which directly (Wang & Isola, 2020; Xu & Durrett, 2018; Wang et al., 2017) and indirectly (Salimans & Kingma, 2016; Franke et al., 2023; Kodryan et al., 2022; Kosson et al., 2023) support representation learning on the hypersphere. One of our main contributions is the normalization of the embedding dimensions of all matrices, ensuring they reside on the same hypersphere. Crucially, we observed that such normalization alone would constrain the inputs of non-linear units, and, thus, the scaling factors for these units should be introduced.
@@ A.2 Eigen learning rates
A.3 Riemannian optimization
@@
A.4 Time cost per step
@@ A.5 Experimental setup
We trained our models using 64 A100 GPUs distributed across 8 nodes (8 GPUs per node). Global batch size is 512. We use the LLaMA-2 tokenizer with 32k tokens. We use the same setup for the 0.5B and 1.0B parameter models.
@@
A.6 Selection of the initial learning
@@
A.7 Length Extrapolation Ability
@@
Figure 13: PG19 perplexity from 1K to 32K.
A.8 Ablation studies
@@
In Table 6, we investigate the impact of removing certain normalization operations in nGPT to potentially reduce training time per step. We remove QK normalization by eliminating the normalization terms Norm in equations 15 and 16. The results show that this modification lead to similar accuracy and loss compared to the baseline, indicating that they are effective options for reducing computational overhead without compromising performance.
@@
Table 6: Ablations on the removal of QK normalization in nGPT which involves taking out the normalization terms Norm in equations 15 and 16.
ARC-E HellaSwag WinoGrande WSC273 LAMBADA Avg. Acc (%) $\uparrow$ Valid. Loss $\downarrow$
Baseline nGPT 53.07 46.03 55.96 67.03 50.13 54.44 2.252
Baseline nGPT - QK norm 52.65 46.07 56.27 68.86 49.68 54.71 +0.12%
A.9 Analysis of scaling parameters

### v2

@@ 1 Introduction
We propose to unify (see also Appendix A.3) various findings and observations made in the field under a new perspective of the normalized Transformer. Our key contributions are as follows:
@@
Faster convergence We demonstrate that the normalized Transformer (nGPT) reduces the number of training steps required to achieve the same accuracy by a factor of 4 to 20.
@@ 2.1 Token Embeddings and Output Logits
Here, $z_{i,y_{i}}$ is the logit corresponding to the correct token $y_{i}$, and the denominator normalizes the logits into a probability distribution over the vocabulary. During inference, the prediction $\hat{{\bm{y}}}_{i}$ is obtained by selecting the token with the highest probability. Since all nGPT embeddings are normalized, the logits ${\bm{z}}\in\mathbb{R}^{V}$ in equation 1 represent dot products bounded in the range $[-1,1]$. This limits the confidence (temperature) of the probability distribution generated by the softmax in equation 2. To adjust this during training, and in line with Hoffer et al. (2018), we introduce a trainable scaling parameter ${\bm{s}}_{z}\in\mathbb{R}^{V}$ that scales the logits element-wise:
@@ 2.6 Summary of modifications
    After each training step (and, optionally, during the forward pass), normalize matrices ${\bm{E}}_{\text{input}}$, ${\bm{E}}_{\text{output}}$, ${\bm{W}}_{q}$, ${\bm{W}}_{k}$, ${\bm{W}}_{v}$, ${\bm{W}}_{o}$, ${\bm{W}}_{u}$, ${\bm{W}}_{\nu}$ and ${\bm{W}}_{o\text{MLP}}$ along their embedding dimension.
@@ 3.1 Acceleration of training
Figure 3 shows a similar pattern across downstream tasks, confirming that the acceleration is not only reflected in perplexity but also in task performance. Figures 8 and 10 in the Appendix provide results for 1k and 8k context lengths. We observe some saturation for the longest runs of nGPT, suggesting that the model capacity is nearly reached for this number of trainable model parameters.
@@ 3.2 Inspection of network parameters
An important contribution of this work is the decoupling of predictions made by the Attention and MLP blocks from their impact on the hidden state ${\bm{h}}$. These contributions are controlled by the eigen learning rates $\bm{\alpha}_{\text{A}}$ and $\bm{\alpha}_{\text{M}}$. Their interpretation is straightforward: if $\bm{\alpha}_{\text{A},i}$ for an embedding dimension $i\in\mathbb{R}^{d_{\text{model}}}$ is 0.2, then the update follows $h_{i}\leftarrow(1-0.2)h_{i}+0.2h_{\text{A},i}$. Thus, they directly quantify the contribution of $h_{\text{A},i}$ into ${\bm{h}}$. Figure 6 shows the average absolute values of ${\bm{h}}_{\text{A}}$ and ${\bm{h}}_{\text{M}}$ at each layer. Notably, the network learns to take only modest steps (20%-30%) in the direction suggested by ${\bm{h}}_{\text{A}}$ and ${\bm{h}}_{\text{M}}$. The average magnitude of $\bm{\alpha}_{\text{A}}$ decreases from 0.25 in the 0.5B network (24 layers) to 0.20 in the 1B network (36 layers). Meanwhile, $\bm{\alpha}_{\text{M}}$ decreases from 0.37 to 0.32. If the parameter count per block is linked to its maximum predictive capacity (in our setup, MLP blocks have more parameters than Attention blocks), then the greater eigen learning values observed for MLPs could potentially be attributed to the higher quality predictions made by these blocks.
@@ 4 Related Work
The normalization in equations 15 and 16 closely resembles the QK normalization by Henry et al. (2020). In nGPT, this process can be viewed as restoring ${\bm{q}}$ and ${\bm{k}}$ of the $i$-th head to a $(d_{\text{model}}/n_{\text{heads}})$-dimensional hypersphere after the projection of ${\bm{h}}$ by ${\bm{W}}_{q}$ and ${\bm{W}}_{k}$, respectively. Since ${\bm{h}}$ and the embedding dimensions of ${\bm{W}}_{q}$ and ${\bm{W}}_{k}$ are already normalized, the norms of ${\bm{q}}$ and ${\bm{k}}$ are also comparable, making their normalization potentially unnecessary. We investigated the effect of omitting this normalization in our ablation studies (see Appendix A.9). The results indicate only a minor performance degradation with potential computational savings of about 12%. However, the normalization helps to maintain performance when extrapolating the context length (see Appendix A.8).
@@ 5 Discussion and Conclusion
This work builds on numerous key findings and observations made in the field which directly (Liu et al., 2017; Wang et al., 2017; Liu et al., 2018; Xu & Durrett, 2018; Wang & Isola, 2020; Liu et al., 2021; Karras et al., 2024) and indirectly (Salimans & Kingma, 2016; Franke et al., 2023; Kodryan et al., 2022; Kosson et al., 2023) support representation learning on the hypersphere. One of our main contributions is the normalization of the embedding dimensions across all of the transformer’s matrices, ensuring they reside on the same hypersphere. Crucially, we observed that such normalization alone would constrain the inputs of non-linear units, and, thus, the scaling factors for these units should be introduced.
@@ A.2 Eigen learning rates
During the review process of this paper, we were made aware of the work of Bachlechner et al. (2020) which proposes ReZero (residual with zero initialization). In the context of Transformers, ReZero is implemented as:
${\bm{h}}$ $\leftarrow{\bm{h}}+\alpha{\bm{h}}_{\text{T}}$ (24)
where the hidden state ${\bm{h}}$ at each layer is updated with $\alpha$-rescaled output ${\bm{h}}_{\text{T}}$ of the transformation block. The trainable layer-specific scalar $\alpha$ is the same for both the MLP and Attention blocks, with its initial value suggested to be set to zero. The authors also proposed to remove LayerNorm and suggested that BatchNorm is a suitable alternative. We note that the introduction of the scaling factor $\alpha$ does not affect the architecture of the model because the output projection matrices of both MLP and Attention blocks already contain that same scaling factor as a multiplicative part of their parameters (e.g., a simultaneous rescaling of $\alpha$ by $k$ and the output projection matrices by $1/k$ results in a network with identical behavior). However, the use of a single scaling factor can help to more quickly adjust the impact of each transformation block. Indeed, the authors argue that by initializing $\alpha$ to zero, one can achieve a form of curriculum learning in which individual layers of the network can be gradually activated during optimization.
The update of ReZero differs from the update of nGPT which can be written as:
${\bm{h}}$ $\leftarrow\text{Norm}(\,{\bm{h}}+\bm{\alpha_{\text{T}}}(\text{Norm}({\bm{h}}_{\text{T}})-{\bm{h}})\,),$ (25)
where the original output of the transformation block ${\bm{h}}_{\text{T}}$ is normalized before being used to compute an update direction, weighted by the eigen learning rate of the block $\bm{\alpha_{\text{T}}}$. In contrast to ReZero, the update of nGPT is invariant to any rescaling of the output projection matrix, and, thus, the eigen learning rate can be viewed as a measure of the contribution of the block. This is not the case for ReZero, where the scale of $\alpha$ cannot be viewed as a measure of the impact of the block without taking into account the scaling of the block’s matrices (see, e.g., our previous example with rescaling by $k$). In contrast to ReZero, where $\alpha$ rescales ${\bm{h}}_{\text{T}}$, the update of nGPT rescales the direction towards ${\bm{h}}_{\text{T}}$. Finally, the update of nGPT holds for the spherical representation where the updated hidden state is retracted back to the hypersphere.
Another work that we were made aware of during the review process is NormFormer (Shleifer et al., 2021) which updates the hidden state as follows:
${\bm{h}}$ $\leftarrow{\bm{h}}+\text{LN}(\,{\bm{h}}_{\text{A}}\,)$ (26)
${\bm{h}}$ $\leftarrow\bm{\alpha}{\bm{h}}+\text{LN}(\,\sigma(\text{LN}({\bm{h}})){\bm{W}}_{1}){\bm{W}}_{2},$ (27)
where ${\bm{h}}_{\text{A}}$ is the output of the Attention block with LayerNorm normalization of inputs, ${\bm{W}}_{1}$ and ${\bm{W}}_{2}$ matrices are components of the MLP block with GELU activation function, and, $\bm{\alpha}$ are trainable scaling factors defined parameter-wise. If we omit the variance part of LayerNorm in Equation 26, then $\text{LN}(\,{\bm{h}}_{\text{A}}\,)$ can be reshaped to $\bm{\alpha}\text{Norm}(\,{\bm{h}}_{\text{A}}\,)$ and one could wonder if $\bm{\alpha}$ can be viewed as eigen learning rates. This is not the case because, similarly to ReZero, i) the norm of ${\bm{h}}$ is not constrained, and ii) the scaling is applied to the value itself and not to the direction. In Equation 27, not only the norm of ${\bm{h}}$ is not constrained, but the output of the MLP is also not normalized. Instead, LayerNorm normalizes the rescaling factors for the vectors stored in ${\bm{W}}_{2}$. While the name NormFormer may suggest some similarities with nGPT, the approach does not normalize network parameters in any way and does not normalize embeddings during the update over layers (i.e., the typically growing norm of ${\bm{h}}$ affects its recombination with the outputs of the transformation blocks).
A.3 Supplementary Notes to Introduction
The original Transformer architecture is a solution in some design space which can be measured with respect to a set of metrics. Many known aspects of multi-criteria decision-making and multi-objective optimization are applied (Deb et al., 2016). Various Transformer variants exist because depending on the selection of metrics (e.g., performance on particular tasks) and constraints (e.g., memory, compute, data), the decision-maker prefers and selects (e.g., based on non-dominance or aggregation of objectives) different design solutions. The first part of the Introduction section describes various architectural and optimization findings/solutions of the past. Then, it is proposed to ”unify various findings and observations made in the field under a new perspective of the normalized Transformer.” This unification can be viewed as a recombination procedure in the design space of architectures and optimization approaches. The Introduction section concludes with a description of the key aspects of the proposed solution. The experimental results of the paper demonstrate some advantages of this solution w.r.t. the baseline GPT, e.g., improved condition number of attention matrices, faster convergence.
While the proposed solution can be viewed as a recombination of existing techniques, this is not how it was historically designed. The precursor of nGPT is described in Loshchilov (2023), where a modification of AdamW was studied in the training of Transformer models. That work demonstrated that instead of performing weight decay, the norm of all parameters of the network can be controlled and scheduled over the course of training. While this approach enabled acceleration in training, it was noted that the attention matrices are ill-conditioned and the normalization of the whole network or even individual matrices does not resolve it. In contrast, experiments with normalization of individual vectors within matrices demonstrated improvements of the condition numbers. Then, by combining this finding with the understanding that LayerNorm and RMSNorm normalize the hidden state to a hypersphere (when all scaling factors are equal), it was hypothesized that the spherical representation is a viable option. nGPT was designed from first principles, focusing on ensuring that the Transformer architecture respects the spherical representation for all its vector components. SLERP was identified as the proper way to recombine the hidden state and the output of the transformation blocks when working on the hypersphere and using scalar eigen learning rates. Scaling factors, such as those used for logits, were introduced to relax the constraints imposed by the spherical formulation.
During the review process of this paper, we were made aware of a set of closely related works. Karras et al. (2024) introduces a set of normalization techniques (very similar to ours) which greatly improve the training of Diffusion Models. Liu et al. (2021) proposed an optimization method for neural networks under hard hyperspherical weight constraints. Large et al. (2025) employs a residual block structure which is related to the one of nGPT but without learnable parameters (eigen-learning rates). By connecting transformers to modern Hopfield models, Wu et al. (2024) show that well-separated memory patterns (learned representations) enhance memorization capability (expressiveness), naturally achieved by mapping learning onto the hypersphere. The authors conclude that representation learning on the unit hypersphere significantly improves expressiveness and memorization. Hu et al. (2024) claim that mapping to the hypersphere makes representation learning provably optimal for any dataset.
A.4 Riemannian optimization
@@
A.5 Time cost per step
@@ A.6 Experimental setup
We trained our models using 64 A100 GPUs distributed across 8 nodes (8 GPUs per node). Global batch size is 512. We use the LLaMA-2 tokenizer with 32k tokens. We use the same setup for the 0.5B and 1.0B parameter models. All parameters of matrices are stored in bfloat16.
@@
All experiments described in this paper were performed using an internal library based on Megatron-LM (Shoeybi et al., 2019). In order to illustrate how nGPT works, we re-implemented nGPT using nanoGPT (Karpathy, 2023) and published our re-implementation at https://github.com/NVIDIA/ngpt. It should be noted that the latter re-implementation only qualitatevely replicates our internal experiments and should not be used in production. However, it should be great to quickly get familiar with nGPT as its core functionality represents only a few dozens of lines of codes. It is very important to note that when implementing nGPT in training libraries, one should make sure that not only instantiated model parameters are normalized but also the ones which are used by the optimizer. Missing the latter is a common bug that should be avoided.
A.7 Selection of the initial learning
@@
Figure 10: Performance on WMT14-FR-EN task which evaluates the ability (measured by the BLEU score of Papineni et al. (2002)) to translate a French sentence to English given five (French, English) example pairs, following the methodology of Radford et al. (2018) where it was used to benchmark GPT-2.
@@
A.8 Length Extrapolation Ability
@@
In the original version of the paper, we explored the impact of qk-normalization on performance of nGPT. Our conclusion was that the performance is equivalent but removing qk-normalization can provide computational savings. During the review process of the paper, two reviewers asked questions which made us further investigate the role of qk-normalization and context length extrapolation abilities of nGPT. This resulted in noticing that nGPT without qk-normalization shows worse extrapolation abilities (see Figure 14-Right). This difference was initially overlooked because the results of nGPT with and without qk-normalization on other tasks are practically equivalent (see, e.g., the average accuracy on downstream tasks and validation loss of both variants in Table 6). In other words, their in-distribution performance appears the same (i.e., qk-normalization does not explain nGPT’s better performance) but their out-of-distribution performance is different. The difference comes from the fact that qk-normalization normalizes ${\bm{q}}$ and ${\bm{k}}$, and, thus, guarantees that their dot product is bounded in [-1,1]. Even without using qk-normalization in nGPT, the hidden state and the vectors forming ${\bm{W}}_{q}$ and ${\bm{W}}_{k}$ matrices are normalized so that their dot product is already bounded in [-1,1]. However, the results of these individual dot products (based on vectors from ${\bm{W}}_{q}$ and ${\bm{W}}_{k}$) will form ${\bm{q}}$ and ${\bm{k}}$ per head. Moreover, ${\bm{q}}$ and ${\bm{k}}$ are also affected by RoPE. The qk-normalization restores ${\bm{q}}$ and ${\bm{k}}$ vectors back to a hyper-sphere whose dimensionality is defined by the size of each head.
While the extrapolation abilities of nGPT can be partially attested to the use of qk-normalization, the gap between nGPT without qk-normalization and the baseline GPT is still substantial. Another contributing factor to the difference is the fact that many attention matrices of the baseline GPT are practically low-rank, as discussed in the main paper. This makes them less efficient in dealing with the context. Interestingly, Kobayashi et al. (2024) recently demonstrated that weight decay induces low-rank attention layers, thus confirming our observations. While the low-rank attention matrices can potentially appear due to a set of factors, we note that weight decay is not used in nGPT.
Figure 14: PG19 perplexity from 1K to 32K among different training lengths for the original GPT and nGPT (Left) and also with nGPT when qk-normalization is removed (Right).
A.9 Ablation studies
@@
In Table 6, we investigate the impact of removing QK normalization operations and the approximation of using LERP instead of SLERP to potentially reduce training time per step. Specifically, removing the QK normalization terms (Norm) in equations 15 and 16 reduces training time per step by 12%. Conversely, replacing the baseline LERP in equation 7 with SLERP in equation 6 increases training time by 10%. Despite these changes, both modifications result in comparable accuracy and loss to the baseline, demonstrating their effectiveness in reducing computational overhead without sacrificing performance.
@@
Table 6: Ablations of the design choices of nGPT. The removal of QK norm involves taking out the normalization terms Norm in equations 15 and 16. The replace of LERP with SLERP is using equation 6 without approximation of equation 7.
Training time per step (s) $\downarrow$ Avg. Acc (%) $\uparrow$ Valid. Loss $\downarrow$
Baseline nGPT 0.657 54.44 2.252
1) remove QKnorm 0.576 54.71 +0.12%
2) replace LERP with SLERP 0.726 54.80 -0.08%
A.10 Analysis of scaling parameters